Part C & D please

2. (30 points) Consider an investment in units of health capital that follows a function: H500-(800 cost of capital). Suppose that Sam starts his year with 5,000 units of health capital. (a) (5 points) If the interest rate is 4% per year and that the depreciation rate is 6% per year Assumung that Sams health does not appreciate in value, what is his cost of capital? (b) (10 points) If Sam were to engage in no health investment this year, what would happen to his capital (c) (5 points) If Sam does invest in health capital what will be his gross (total) investment, in terms of units (d) (10 points) Assuming that the interest rate were to rise to 6% per year in the following year, calculate stock? of capital? what would happen to the equilibrium level of investment units of health capital.
Answer #1

c) 5000(1.04)=5200 => now incorporating depreciation 5200=500-(800*C)

Gross investment =-5.875

d) 5200*0.06=500-(800*Y)

Y= 0.235